# Numerical efficiency of explicit time integrators for phase-field models

### v1

@@ 4.3 Single grain growth
In this case only the phase-field equation eq. 6 is solved for, as both grains are of the same phase. The inner grain uses the distance field eq. 41, with the outer one being the complement. The area of the grain follows from simple volume integration.
4.4 Solutal Stefan problem
The second kinetic test approximates the classical Stefan problem in a solutal setting. Two semi-infinite phases $\alpha$ and $\beta$ are brought into contact with respective initial compositions $c_{\alpha}$, $c_{\beta}$. If these differ from their equilibrium composition, then the interface between them will move. Following Balluffi et al. [2005], employing a similarity solution for the diffusion fields and incorporating the flux balance at the interface as well as assuming equal atomic volumes and diffusivities $D$, yields
@@ 4.4 Solutal Stefan problem
with suitably defined Gibbs-Thomson coefficient $\Gamma$ and kinetic coefficient $K$. Employing a one-dimensional geometry ensures $\kappa$ vanishes. Since $v>0$ is required for kinetics, we are left with making $K$ vanish by determining its relation to the phase-field parameters. The latter is usually done with a thin-interface analysis; for details, refer to e.g. Choudhury and Nestler [2012]. A relationship for the phase-field mobility $L$
@@
The distance field for phase $\beta$ is given by d=hxd=h-x for an initial height $h$ from the left boundary, with that of $\alpha$ being the complement. The initial concentration is interpolated as $c(t=0)=\phi_{\alpha}c_{\alpha}+\phi_{\beta}c_{\beta}$, i.e. the far-field phase concentration values are employed. For the phase-field, gradient zero boundary conditions are employed, whereas the boundary concentrations are held at the respective initial concentrations with Dirichlet conditions.
@@
We note that there is an additional error source here compared to the other benchmark geometries, namely the far-field concentration. For the analytical solution an infinite domain is assumed, which obviously cannot be simulated. Hence one must ensure that the gradients in the far field are negligible. We chose a time giving a certain displacement $X(t)=50\to t\approx$43\text{\times}{10}^{3}$\$ via eq. 50 and then varied the physical domain size in an informal study to fix it for the reported simulations. Finally, the momentary position $h+X(T)$ is determined based on the position of the $\phi_{\alpha}=0.5$ point determined via solving a cubic interpolant.
4.5 Grain growth with particles
As an application example we will also treat a system which doesn’t have an analytical solution, but which is representative of more typical ratios of kinetic parameters as well as more complicated geometries. For this a system of many grains $\beta$ of different orientation is employed, together with spheres of a second phase $\alpha$ randomly placed in this structure. When these second phases are in contact with a grain boundary, they exert a drag force on the boundary. Depending on the effective mobility of the second phase, these may be taken to be immobile particles which act as pinning centers, allowing for investigations relating to Zener pinning and limiting grain sizes. For more mobile phases, this represents the final stage of sintering with only closed porosity remaining, which reduces grain growth but does not necessarily lead to a limiting grain size, as pores may move and end up merging. The effective particle mobility can be controlled via the size of the second phase and the diffusivity controlling the mass transportLiu and Patterson [1994], which for simplicity here was taken to be diffusion within the second phase, i.e. vapour diffusion if we interpret it as a pore. We note that Zener pinning can in general be more easily simulated without recourse to using a concentration field, see e.g. Harun et al. [2006]. The final sintering stage can also be approximated without diffusion fields, see e.g. Rehn et al. [2019], but this assumes known dominant diffusion mechanisms, more materials parameters and a tracking algorithm to account for pore pressure and pore-pore interactions. Furthermore, pores in the final sintering stage can also be elongated, in which case diffusional instabilities such as the Plateau-Rayleigh instability become relevant. These are more easily described by employing concentration fields. We note that the eventually chosen parametrization for this case may be calculated more efficiently by integrating the phase-field and concentration fields separately, as their respective timescales differ by several orders of magnitude; but for ease of presentation and comparability both fields were integrated together.
We seek to reproduce the following qualitative behaviours with three simulations: If grain growth without particles happens, mean-field theory based on curvature as the sole driving force predicts parabolic growth of the mean grain size, i.e. $G(t)^{n}-G(0)^{n}=kt,n=2$Rahaman [2017]. As alluded to in the previous paragraph, immobile particles can lead to a limiting grain size, which in experiments may also be identified as large exponents $n$ since the curve will tend to flatten out increasingly. In contrast to this, while mobile particles (pores) also exert a drag force on the grain boundary, they co-move with the grain boundary and may merge with others, reducing their effectiveness as pinning centers. Hence in this case the exponent generally increases, see e.g. Nichols [1966], Riedel and Svoboda [1993], while not necessarily reaching a limiting grain size. Based on the different parameters here we will be able to observe speedups for realistic parameter combinations.
The initial conditions for pure grain growth are a periodic Voronoi tesselation of 3D space with random points, with a minimal distance of $12\Delta x$ between each point of a $(400\Delta x)^{3}$ domain. Once this simulation has reached a roughly parabolic regime ($t=1$ in the computational experiments) simulations with particles are constructed based on this configuration: For immobile particles, spheres are placed randomly in space with a target volume ratio $f_{i}=0.15$ and radius $r_{i}=6\Delta x$ in order to showcase the limiting grain size, with the distance between spheres being at least $d=2r$. Due to the finite extent of the phase-field this still means that some overlap exists, which does increase the effective radius but due to the low mobility the particles do not move or change shape over the simulation duration. However, for mobile particles this would lead to quick merging of most particles, giving a completely different geometry than expected. Hence for these a minimal distance of $d=2(r+e),e=8\Delta x$ is imposed, which limits the obtainable porosity as a function of this distance and the pore radius. Due to this limit we pick $f_{m}=0.05,r_{m}=10\Delta x$ which is also more representative of the porosity in final stage sintering. While it is possible to also calculate the immobile particle case with this particle distribution, the simulation domain would need to be enlarged significantly to reach the limiting grain size. For the immobile parameters the classical Zener pinning theory predicts a limiting grain size of $r_{l,i}=\frac{4}{3}\frac{r}{f}\approx 53.3\Delta x$ whereas for the mobile particle distribution it would be $r_{l,m}\approx 266\Delta x$ and hence require a domain in excess of $(800\Delta x)^{3}$Miodownik et al. [2000]. While calculating such a domain poses no problems but the waiting time for the employed code, the goal of reproducing qualitative behaviour does not necessarily justify the simulative expense.
We note that the phase-field volume $\int_{V}\phi(d)dV$ of a two and higher dimensional shape has an $O(W^{2})$ error compared to the sharp interface volume the distance field $d$ is supposed to represent. In order to still obtain a porosity close to the target porosity even when the size of the particles is comparable to the interface width, the radius $r$ is transformed to a smaller equivalent radius $r_{c}$ such that the target volume is obtained with the help of a function function $r_{c}=f(r,W)$ derived in the supplementary material. With the centers and radius of the particles fixed, the only thing left is to place them in the domain. Since a diffuse initial shape is used, $\phi_{\alpha}=\phi(r_{c}-x_{min})$ is imposed for each cell while keeping the ratios of the remaining phase-fields constant as detailed in section 4.2. $x_{min}$ here is the smallest distance of this cell to any of the particle positions. The geometry is hence fully described and what remains is to fix the parameters representing the different regimes.
For insoluble, mobile particles (pores) we want small particle in some sense with sufficient diffusivity. Since the mobility of these scales as $r^{-n},n>2$ Liu and Patterson [1994] with the particle size $r$, choosing a smaller length scale means we have more mobile particles relative to pure grain growth. At the same time, we would like to use the same length scale for immobile particles, hence after fixing the length scale, diffusion values leading to immobile and mobile particles are determined. Furthermore, in order for the particles to be insoluble even under capillary action the prefactor $k$ in eq. 17 needs to be large relative to capillary forces. We can estimate this effect by considering the expression for the phase concentration in terms of the chemical potential $c_{\alpha}=c_{0,\alpha}+\frac{\mu}{k}$ and the equilibrium condition $\psi_{\alpha}-\psi_{\beta}=\gamma\kappa$, from which we find that the shifted equilibrium bulk concentration is $c_{\alpha}(\gamma\kappa)=c_{0,\alpha}-\frac{\gamma\kappa}{k(c_{0,\alpha}-c_{0,\beta})}$. We would like for the difference to the bulk concentration to be small relative to the bulk concentration, and hence must know the scale of $O(\kappa)=O(\frac{1}{r})$. By computational experiments we pick a length scale of $\Delta x=0.005$ at which $D_{\alpha}=10$ shows visibly mobile particles and hence $k=$1\text{\times}{10}^{5}$\$ ensures $\frac{c_{\alpha}(\gamma\kappa)-c_{0,\alpha}}{c_{0,\alpha}}<1\%$ for the considered particle sizes.
If the immobile particle parameters are employed directly after setting the particles, stable many phase regions tend to form on the fringes of the particles, even if a simplex accounting for mobility variations is employed as suggested in Daubner et al. [2023]. We suspect that this is due to chemical driving forces aiming to shrink the particles to achieve their capillary equilibrium, which gives a driving force to any grain phases in the vicinity. This driving force may exceed the potential energy terms of the phase-field and hence stabilize many phase regions which originally exist due to the interface width being comparable to some of the initial grains’ equivalent radius. While these would eventually vanish after capillary equilibrium is achieved, the timescale for that is far beyond the timescale of grain growth for this parameter combination. Since these may act as artificial pinning centers, an additional equilibration run is conducted for both particle-laden simulations prior to running with the actual parameters. The run uses $M_{\beta\beta}=$5\text{\times}{10}^{-8}$,D_{\alpha}=$5\text{\times}{10}^{-3}$,M_{\alpha\beta}=$5\text{\times}{10}^{-3}$\$ as to ensure that virtually no grain growth happens while at same time allowing for capillary relaxation, and is continued for a duration of $5$ time units. After this equilibration run the time is reset to $t=1$ and the simulations are run normally; no further formation of stable many phases regions was observed.
5 Results
In this section the results of the simulations studies are presented. We first ascertain whether the STS time integrator approximates the sharp interface limit for equilibrium configurations. This is followed by determining whether the kinetics are matched as well, together with a performance comparison between probed integrators. Performance is reported here in terms of RHS evaluations, as this measure is independent of the machine employed and hence eases future comparison with other time integrators. Qualitative comparison to implicit integrators can also be done with this as long as the resulting systems are solved with iterative methods as each iteration requires RHS evaluations; a more quantitative comparison needs to account for e.g. additional matrix-vector operations required by the solution algorithm as well as potential mass matrix inversions. The supplementary material includes data on the runtimes as well for the interested reader, though the conclusions do not change if performance is measured in runtime instead of RHS evaluations.
The main set of integrators with their configurations is described in Table 2. If not mentioned otherwise, adaptive integrators use tolerances of $a=$1\text{\times}{10}^{-4}$,r=$1\text{\times}{10}^{-4}$\$ for all fields. For grid convergence studies, only the STS2 integrator with $a_{\phi}\in\{$1\text{\times}{10}^{-2}$,$1\text{\times}{10}^{-4}$\$\} is employed to reduce computational expense. The SSP(10)4 integrator at tolerances of $1\text{\times}{10}^{-14}$ is generally used as a reference solution generator for testing convergence order.
Table 2: Employed integrator configurations
integrator configurations
FEuler $\Delta t\in\{0.5,1\}\cdot\Delta t_{e}$
SSP(5)2 $\Delta t\in\{0.5,1,2,4\}\cdot\Delta t_{e}$
STS[12]
$\Delta t\in\{1,10,100,200\}\cdot\Delta t_{e}$
Finally, we generally output at defined target times exactly. For fixed-step integrators this only manifests in an extra RHS evaluation for each output time not being an integer multiple of the time step width. For adaptive integrators, this generally reduces the step taken prior to reaching the output significantly and hence becomes less efficient than if a simple time crossing output strategy were adopted. Hence the performance results are slightly biased against adaptive integrators.
5.1 Equilibrium verification
First, we perform a mesh convergence study with the geometry of section 4.1, resolving a domain of $128^{2}$ with an initial seed of size $r=32$. Equilibrium for this case is defined as a rate of change for the Laplace pressure $|\frac{(\Delta p(t)-\Delta p(t_{b}))}{t-t_{b}}|$ being smaller than $1\text{\times}{10}^{-14}$ for two times $t>t_{b}$ at least two diffusion times $128^{2}/D$ apart.
(a) errors in the Laplace pressure
(b) energy profile over time
For the triple junction geometry, a domain of size $192\times 96$ is resolved at the coarsest scale with $\Delta x=1$, with the long side being along the grain boundary to facilitate simulation of smaller dihedral angles at the same aspect ratio. The initial radius of the spherical phase on the grain boundary is $32$. For simplicity the convergence study is limited to a single angle $\gamma_{\beta\beta}=1,\gamma_{\alpha\beta}=2\rightarrow\theta\approx$151\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, as results should be transferable to other angles except when approaching $\theta\to 0$ as complete dewetting is not captured by the model Daubner et al. [2023]. In addition, the coarsest resolution is also run with the FEuler integrator with a time step 0.9 times its stable time step. The remaining angles in the range $95\text{\,}\mathrm{\SIUnitSymbolDegree}180\text{\,}\mathrm{\SIUnitSymbolDegree}$ will use the practical resolution $\{\Delta x=1,W=2.5\}$ with $\gamma_{\alpha\beta}\in\{0.75,1,2,6\}$.
Equilibrium is defined as a simple approximation to the measurement described in section 4.2. The positions of the 0.5 lines are determined on the horizontal and vertical lines passing through the center of the domain, using only a linear interpolation in the respective coordinate directions, with their respective differences giving approximations to $L,S$. This allows an estimate of the current dihedral angle $\theta$, with equilibrium defined as a rate of change in angle by less than $1\text{\times}{10}^{-11}\text{\,}\mathrm{rad}$ over at least two diffusion times. The reported angle values are based on a more accurate marching squares reconstruction of the 0.5 contour lines of the phase $\alpha$, which is only done after the simulation is completed.
(a) errors in the dihedral angle
(b) energy profile over time
(c) time evolution of the angle
The results of the grid convergence study are shown in fig. 3. Again, convergence to the sharp interface is observed with little influence from $a_{\phi}$. At the coarsest resolution, a relative error of $1.4\%$ is obtained. In this case, the energy does not decrease monotonically beyond the first output frame as shown in the inset of fig. 3(b); while there is a long-term decrease of energy, it is not monotonic. As is visible from the graph, on the scale of the energy reduction it is hardly noticeable without zooming closely to the data. Note that the FEuler integrator at the coarsest resolution also exhibits an non-monotonic behaviour as shown in the top inset. Rejection rates ranged from $1\%$ for $a_{\phi}=$1\text{\times}{10}^{-2}$\$ to up to $27\%$ for $a_{\phi}=$1\text{\times}{10}^{-4}$\$. It is unlikely that step rejection is influenced by energy instability due to this split. We note though that energy instability was only apparent once the energy was almost converged; benchmarks not getting close to equilibrium are unlikely to probe energy stability properties of the integrator.
[…]
5.2 Kinetic verification
Now that we know that the lack of strict energy stability does not significantly impact the equilibrium we can proceed to probe the kinetics of the integrators and get an impression of their performance. First is the embedded grain geometry described in section 4.3. We start again with a mesh convergence study to show that the proper limits are recovered, with a domain of size $512^{2}$ being employed and an initial radius of $r=220$ for the embedded grain, with the simulation being terminated upon reaching $t=$23\text{\times}{10}^{3}$\$ before the grain vanishes.
(a) area evolution
(b) spatial convergence
(c) temporal convergence for $W=2.5$
(d) performance for $W=2.5$
The results of this study and one using the integrators specified in Table 2 are shown in fig. 4. In fig. 4(a), we can observe that the qualitative time dependence of the exact solution is obtained by all simulations, but with quantitative differences. In fig. 4(b) convergence is observed throughout for time integration tolerances of $1\text{\times}{10}^{-4}$, but for $a_{\phi}=$1\text{\times}{10}^{-2}$\$ this eventually limits the convergence. Interestingly, the error is smaller for the relaxed tolerance prior to this, suggesting that lower error may be achievable using a coarser tolerance due to cancellation of spatial and temporal errors. Thus for practical simulations, using a relatively coarse absolute tolerance can be beneficial for both speed and accuracy. Outside of the convergence part, the effect of time integration error is generally small as evidenced by the close clustering of crosses representing the remaining simulations. The largest of these, corresponding to a STS1 scheme with the largest timestep, has a relative error of about $3\%$, with most other simulations clustering around $2.7\%$ relative error.
@@ 5.2 Kinetic verification
The next case employs the Stefan problem as described in section 4.4. A one-dimensional domain of size 1800 is employed, with the left boundary held at the equilibrium concentration of $\beta,c_{\beta}=0.98$, with the right one held at $c_{\alpha}=0.2$. Together with the diffusivity $D=1$ this results in a growth constant $A\approx 0.2412$ via eq. 50. The initial height of the $\beta$ phase is 400. All results are collected in fig. 5(a) which shows a close match to the exact solution. We observe that the model converges upon grid refinement from fig. 5(b), with the practical resolution results still showing good accuracy. All results appear insensitive to the choice of $a_{\phi}$; informal variation of the remaining tolerances also showed little influence, suggesting the error is mainly controlled by diffuse interface effects, which naturally get scaled down due to the choice of $W(\Delta x)$. Considering that the rate of change of all field variables is much slower than in the previous case, the lack of sensitivity of the time integration error is quite sensible.
@@
For the practical interface width the work-precision chart is shown in fig. 5(c). We observe, as would be suggested by previous results, that the obtained error is largely independent of the integrator and tolerance choice. Using either of the adaptive STS schemes here gives a speedup of about 4-5 with effectively no additional error; fixed timestepping produces speedups from 6 to 10. The phase-field tolerance seems to play little role for the performance here, suggesting that the temporal error is dominated by the concentration tolerance which was fixed for this study. Setting $a_{c}=a_{\phi}$ showed a performance improvement (up to a factor of 11) at no significant change in error.
@@ 5.3 Grain growth with particles
The simulations are performed as described in section 4.5, yielding the grain size curves in fig. 6. For these, tolerances $a_{\phi}=$1\text{\times}{10}^{-2}$,a_{c}=$1\text{\times}{10}^{-3}$\$ were chosen since early tests showed it had little influence as in the Stefan benchmark. An exponent close to $n=2$ is observed for pure grain growth, whereas an increased exponent is observed for grain growth with mobile particles, as would be expected. Theoretically one would expect $n=3$ for mass transport through the vapor with the pore pressure being maintained at capillary equilibrium $p=\frac{2\gamma}{r}$Nichols [1966]. We suspect that this is due transient behaviour between the pure and pore-dragged behaviour, as the initial part of the curves are quite similar; reaching $n=3$ likely requires significantly larger domains to allow for the dynamics to emerge. For the immobile particles we can compare the obtained limiting grain size (constant for about a time span of 200) to the prediction of Zener, which comes out to be $r_{l}\approx 0.27$. The simulation predicts a larger value of about $0.3$, which is likely due to the interface width being of the same order of magnitude as the radius of curvature and overlap between particles effectively increasing the average radius. Hence we can say that even at quite coarse tolerances the STS schemes still recovers the essential features of the free-boundary problem we are concerned with.

### v2

@@ 5.2.1 Single grain growth
The inner grain uses the distance field eq. 44, with the outer one being the complement. The area of the grain follows from simple volume integration. The simulation is terminated upon reaching $t=$23\text{\times}{10}^{3}$\$ before the grain vanishes and the curvature becomes comparable to the interface width.
(a) area evolution
(b) spatial convergence
(c) temporal convergence for $W=2.5$
(d) performance for $W=2.5$
A domain of size $512^{2}$ is employed with an initial seed size of $r=220$ for the embedded grain. The results of applying the integrators specified in Table 2 together with the resolution study are collected in fig. 4. In fig. 4(a), we can observe that the qualitative time dependence of the exact solution is obtained by all simulations, but with quantitative differences. In fig. 4(b) convergence is observed throughout for time integration tolerances of $1\text{\times}{10}^{-4}$, but for $a_{\phi}=$1\text{\times}{10}^{-2}$\$ this eventually limits the convergence. Interestingly, the error is smaller for the relaxed tolerance prior to this, suggesting that lower error may be achievable using a coarser tolerance due to cancellation of spatial and temporal errors. Thus for practical simulations, using a relatively coarse absolute tolerance can be beneficial for both speed and accuracy. Outside of the convergence part, the effect of time integration error is generally small as evidenced by the close clustering of crosses representing the remaining simulations. The largest of these, corresponding to a STS1 scheme with the largest timestep, has a relative error of about $3\%$, with most other simulations clustering around $2.7\%$ relative error.
For grain growth with a single interface, the reported speedups should be universal, as the reduced mobility $\gamma M_{\alpha\beta}$ controls the time scale of the problem and the maximal eigenvalue, and hence changing it is only a scaling of time. Once there are multiple interfaces, as later shown for grain growth, we can expect the speedup for adaptive integrators to decrease, as the size of regions with nontrivial error grows and hence smaller steps must be taken; though at the same time the error committed by the FEuler integrator without error control will grow. If these interfaces have different mobilities, the speedup would increase however as the increased stiffness can be handled by the STS integrator at lower cost than by the FEuler integrator.
5.2.2 Solutal Stefan problem
The second kinetic test approximates the classical Stefan problem in a solutal setting, and hence both eq. 6 and eq. 17 will be used. Two semi-infinite phases $\alpha$ and $\beta$ are brought into contact with respective initial compositions $c_{\alpha}$, $c_{\beta}$. If these differ from their equilibrium composition, then the interface between them will move. Following Balluffi2005 , employing a similarity solution for the diffusion fields and incorporating the flux balance at the interface as well as assuming equal atomic volumes and diffusivities $D$, yields
@@ 5.2.2 Solutal Stefan problem
with suitably defined Gibbs-Thomson coefficient $\Gamma$ and kinetic coefficient $K$. Employing a one-dimensional geometry ensures $\kappa$ vanishes. Since $v>0$ is required for kinetics, we are left with making $K$ vanish by determining its relation to the phase-field parameters. The latter is usually done with a thin-interface analysis; for details, refer to e.g. choudhury2012 . A relationship for the phase-field mobility $L$
@@
The distance field for phase $\beta$ is given by d=hxd=h-x for an initial height $h$ from the left boundary, with that of $\alpha$ being the complement. The initial concentration is interpolated as $c(t=0)=\phi_{\alpha}c_{\alpha}+\phi_{\beta}c_{\beta}$, i.e. the far-field phase concentration values are employed. For the phase-field, gradient zero boundary conditions are employed, whereas the boundary concentrations are held at the respective initial concentrations with Dirichlet conditions. The one-dimensional domain’s left boundary is held at the equilibrium concentration of $\beta,c_{\beta}=0.98$, with the right boundary held at $c_{\alpha}=0.2$. Together with the diffusivity $D=1$ this results in a growth constant $A\approx 0.2412$ via eq. 53. The initial height $h$ of the $\beta$ phase is 400. The momentary position $h+X(T)$ is determined based on the position of the $\phi_{\alpha}=0.5$ point determined via solving a cubic interpolant.
@@
We note that there is an additional error source here compared to the other benchmark geometries, namely the far-field concentration. For the analytical solution an infinite domain is assumed, which obviously cannot be simulated. Hence one must ensure that the gradients in the far field are negligible. We chose a time giving a certain displacement $X(t)=50\to t\approx$43\text{\times}{10}^{3}$\$ via eq. 53 and then varied the physical domain size in an informal study to fix it for the reported simulations, with this ending up being a domain of size 1800.
@@
All results are collected in fig. 5, in which fig. 5(a) shows a close match to the exact solution. We observe that the model converges upon grid refinement from fig. 5(b), with the practical resolution results still showing good accuracy. All results appear insensitive to the choice of $a_{\phi}$; informal variation of the remaining tolerances also showed little influence, suggesting the error is mainly controlled by diffuse interface effects, which naturally get scaled down due to the choice of $W(\Delta x)$. Considering that the rate of change of all field variables is much slower than in the previous case, the lack of sensitivity of the time integration error is quite sensible.
@@
For the practical interface width the work-precision chart is shown in fig. 5(c). We observe, as would be suggested by previous results, that the obtained error is largely independent of the integrator and tolerance choice. Using either of the adaptive STS schemes here gives a speedup of about 4-5 with effectively no additional error; fixed timestepping produces speedups from 6 to 10. The phase-field tolerance seems to play little role for the performance here, suggesting that the temporal error is dominated by the concentration tolerance which was fixed for this study. Setting $a_{c}=a_{\phi}$ showed a performance improvement (up to a factor of 11) at no significant change in error, as shown in the supplementary material.
The speedups reported here depend on the Gibbs free energies by relating to the mobility eq. 56 but also on the diffusivities if these are allowed to vary between the phases, though in this case a quantitative phase-field simulation requires the addition of some model term counteracting the interface width dependent excess solute trapping. If this term introduces extra stiffness, the resulting speedups would again be larger, given that the time scale of the problem and hence the magnitude of its time derivatives are controlled by the sharp interface problem independent of the phase-field model. Furthermore, a larger diffusional contrast is beneficial for the speedup as well, since the error in the low-diffusion regions will be smaller and hence an overall larger timestep may be used to achieve the same error.
@@ 5.3 Grain growth with particles
As an application example we will also treat a system which doesn’t have an analytical solution, but which is representative of more typical ratios of kinetic parameters as well as more complicated geometries. For this a system of many grains $\beta$ of different orientation is employed, together with spheres of a second phase $\alpha$ randomly placed in this structure. When these second phases are in contact with a grain boundary, they exert a drag force on the boundary. Depending on the effective mobility of the second phase, these may be taken to be immobile particles which act as pinning centers, allowing for investigations relating to Zener pinning and limiting grain sizes. For more mobile phases, this represents the final stage of sintering with only closed porosity remaining, which reduces grain growth but does not necessarily lead to a limiting grain size, as pores may move and end up merging, becoming less effective pinning centers. The effective particle mobility can be controlled via the size of the second phase and the diffusivity controlling the mass transportLiu1994 , which for simplicity here was taken to be diffusion within the second phase, i.e. vapour diffusion if we interpret it as a pore. We note that Zener pinning can in general be more easily simulated without recourse to using a concentration field, see e.g. Harun2006 . The final sintering stage can also be approximated without diffusion fields, see e.g. Rehn2019 , but this assumes known dominant diffusion mechanisms, more materials parameters and a tracking algorithm to account for pore pressure and pore-pore interactions. Furthermore, pores in the final sintering stage can also be elongated, in which case diffusional instabilities such as the Plateau-Rayleigh instability become relevant. These are more easily described by employing concentration fields. We note that the eventually chosen parametrization for this case may be calculated more efficiently by integrating the phase-field and concentration fields separately, as their respective timescales differ by several orders of magnitude; but for ease of presentation and comparability both fields were integrated together.
We seek to reproduce the following qualitative behaviours with three simulations: If grain growth without particles happens, mean-field theory based on curvature as the sole driving force predicts parabolic growth of the mean grain size, i.e. $G(t)^{n}-G(0)^{n}=kt,n=2$Rahaman2017 . As alluded to in the previous paragraph, immobile particles can lead to a limiting grain size, which in experiments may also be identified as large exponents $n$ since the curve will tend to flatten out increasingly. In contrast to this, while mobile particles (pores) also exert a drag force on the grain boundary, they co-move with the grain boundary and may merge with others, reducing their effectiveness as pinning centers. Hence in this case the exponent generally increases, see e.g. Nichols1966 , Riedel1993 , while not necessarily reaching a limiting grain size. Based on the different parameters here we will be able to observe speedups for realistic parameter combinations.
The initial conditions for pure grain growth are a periodic Voronoi tesselation of 3D space with random points, with a minimal distance of $12\Delta x$ between each point of a $(400\Delta x)^{3}$ domain. Once this simulation has reached a roughly parabolic regime ($t=1$ in the computational experiments) simulations with particles are constructed based on this configuration: For immobile particles, spheres are placed randomly in space with a target volume ratio $f_{i}=0.15$ and radius $r_{i}=6\Delta x$ in order to showcase the limiting grain size, with the distance between spheres being at least $d=2r$. Due to the finite extent of the phase-field this still means that some overlap exists, which does increase the effective radius but due to the low mobility the particles do not move or change shape over the simulation duration. However, for mobile particles this would lead to quick merging of most particles, giving a completely different geometry than expected. Hence for these a minimal distance of $d=2(r+e),e=8\Delta x$ is imposed, which limits the obtainable porosity as a function of this distance and the pore radius. Due to this limit we pick $f_{m}=0.05,r_{m}=10\Delta x$ which is also more representative of the porosity in final stage sintering. While it is possible to also calculate the immobile particle case with this particle distribution, the simulation domain would need to be enlarged significantly to reach the limiting grain size. For the immobile parameters the classical Zener pinning theory predicts a limiting grain size of $r_{l,i}=\frac{4}{3}\frac{r}{f}\approx 53.3\Delta x$ whereas for the mobile particle distribution it would be $r_{l,m}\approx 266\Delta x$ and hence require a domain in excess of $(800\Delta x)^{3}$Miodownik2000 . While calculating such a domain poses no problems but the waiting time for the employed code, the goal of reproducing qualitative behaviour does not necessarily justify the simulative expense.
We note that the phase-field volume $\int_{V}\phi(d)dV$ of a two and higher dimensional shape has an $O(W^{2})$ error compared to the sharp interface volume the distance field $d$ is supposed to represent. In order to still obtain a porosity close to the target porosity even when the size of the particles is comparable to the interface width, the radius $r$ is transformed to a smaller equivalent radius $r_{c}$ such that the target volume is obtained with the help of a function function $r_{c}=f(r,W)$ derived in the supplementary material. With the centers and radius of the particles fixed, the only thing left is to place them in the domain. Since a diffuse initial shape is used, $\phi_{\alpha}=\phi(r_{c}-x_{min})$ is imposed for each cell while using eq. 41 to update the already present phase-fields. $x_{min}$ here is the smallest distance of this cell to any of the particle positions. The geometry is hence fully described and what remains is to fix the parameters representing the different regimes.
For insoluble, mobile particles (pores) we want small particles in some sense with sufficient diffusivity. Since the mobility of these scales as $r^{-n},n>2$ Liu1994 with the particle size $r$, choosing a smaller length scale means we have more mobile particles relative to pure grain growth. At the same time, we would like to use the same length scale for immobile particles, hence after fixing the length scale, diffusion values leading to immobile and mobile particles are determined. Furthermore, in order for the particles to be insoluble even under capillary action the prefactor $k$ in eq. 18 needs to be large relative to capillary forces. We can estimate this effect by considering the expression for the phase concentration in terms of the chemical potential $c_{\alpha}=c_{0,\alpha}+\frac{\mu}{k}$ and the equilibrium condition $\psi_{\alpha}-\psi_{\beta}=\gamma\kappa$, from which we find that the shifted equilibrium bulk concentration is $c_{\alpha}(\gamma\kappa)=c_{0,\alpha}-\frac{\gamma\kappa}{k(c_{0,\alpha}-c_{0,\beta})}$. We would like for the difference to the bulk concentration to be small relative to the bulk concentration, and hence must know the scale of $O(\kappa)=O(\frac{1}{r})$. By computational experiments we pick a length scale of $\Delta x=0.005$ at which $D_{\alpha}=10$ shows visibly mobile particles and hence $k=$1\text{\times}{10}^{5}$\$ ensures $\frac{c_{\alpha}(\gamma\kappa)-c_{0,\alpha}}{c_{0,\alpha}}<1\%$ for the considered particle sizes.
If the immobile particle parameters are employed directly after setting the particles, stable many phase regions tend to form on the fringes of the particles, even if a simplex accounting for mobility variations is employed as suggested in Daubner2023 . We suspect that this is due to chemical driving forces aiming to shrink the particles to achieve their capillary equilibrium, which gives a driving force to any grain phases in the vicinity. This driving force may exceed the potential energy terms of the phase-field and hence stabilize many phase regions which originally exist due to the interface width being comparable to some of the initial grains’ equivalent radius. While these would eventually vanish after capillary equilibrium is achieved, the timescale for that is far beyond the timescale of grain growth for this parameter combination. Since these may act as artificial pinning centers, an additional equilibration run is conducted for both particle-laden simulations prior to running with the actual parameters. The run uses $M_{\beta\beta}=$5\text{\times}{10}^{-8}$,D_{\alpha}=$5\text{\times}{10}^{-3}$,M_{\alpha\beta}=$5\text{\times}{10}^{-3}$\$ as to ensure that virtually no grain growth happens while at same time allowing for capillary relaxation, and is continued for a duration of $5$ time units. After this equilibration run the time is reset to $t=1$ and the simulations are run normally; no further formation of stable many phases regions was observed. All runs used the tolerances $a_{\phi}=$1\text{\times}{10}^{-2}$,a_{c}=$1\text{\times}{10}^{-3}$\$, with the latter being chosen as early tests showed it had little influence as in the Stefan benchmark.
@@
The three simulations, conducted as described above, yield the grain size curves in fig. 6. An exponent close to $n=2$ is observed for pure grain growth, whereas an increased exponent is observed for grain growth with mobile particles, as would be expected. Theoretically one would expect $n=3$ for mass transport through the vapor with the pore pressure being maintained at capillary equilibrium $p=\frac{2\gamma}{r}$Nichols1966 . We suspect that this discrepancy is due transient behaviour between the pure and pore-dragged behaviour, as the initial part of the curves are quite similar; reaching $n=3$ likely requires significantly larger domains to allow for the dynamics to emerge. For the immobile particles we can compare the obtained limiting grain size (constant for about a time span of 200) to the prediction of Zener, which comes out to be $r_{l}\approx 0.27$. The simulation predicts a larger value of about $0.3$, which is likely due to the interface width being of the same order of magnitude as the radius of curvature and overlap between particles effectively increasing the average radius. Hence we can say that even at quite coarse tolerances the STS schemes still recovers the essential features of the free-boundary problem we are concerned with.